OpenStudy (anonymous):

(a+2)(2a-3) i already have the answer but i have one question about it.. soo... 2a-3 * a+2 -------- 4a-6 2a^2-3a -------------- 2a^2+a - 6 <------ the answer MY QUESTION IS how did it become a (+) on 2a^2((((+))))a-6 the one PLUS sign ?!?! someone please explain :(

OpenStudy (texaschic101):

(a + 2)(2a - 3) ---- using the FOIL method (first,outside,inside,last) First -- multiply first terms in each set... a * 2a = 2a^2 Outside -- multiply outside terms in each set...a * -3 = -3a Inside -- multiply inside terms in each set..2 * 2a = 4a Last -- multiply last terms in each set...2 * -3 = -6 put them all together.. 2a^2 - 3a + 4a - 6 -- combine like terms 2a^2 + a - 6 does that help ?

OpenStudy (texaschic101):

it becomes a plus when combining like terms because -3a + 4a = a

OpenStudy (anonymous):

ahhhhhh that explains it :) thanks! i fogot about the FOIL.. thank you!!!!!!!!!

OpenStudy (anonymous):

forgot*

OpenStudy (texaschic101):

anytime :)